OpenStudy (anonymous):

the function described by h(x)=2.75 x +71.48 can be used to predict the height in centimeters of a woman whos humerous (the bone in the elbow to the shoulder) is x cm long predict the height of a woman whose humerus is 33 cm long the predicted height of a woman whose humerus is 33 cm long is how many centimeters

OpenStudy (anonymous):

Question is unclear.

OpenStudy (anonymous):

h(x)=2.75 x +71.48 x=33 cm height of a woman=(2.75*33) +71.48=162.23cm